I need assistance in finding the dimensions. When I work the problem I keep getting a fraction in the answer. How do I find the dimensions?
The perimeter of a rectangular oceanfront lot is 190m. The width is one fourth of the length. Find the dimensions. thank you

Answer by Alan3354(69443)

The perimeter of a rectangular oceanfront lot is 190m. The width is one fourth of the length. Find the dimensions.
----------
Perimeter = 2W + 2L = 190
W + L = 95
L = 4W
W + 4W = 95
5W = 95
W = 19 m
L = 76 m
------------
PS Sometimes the answers are fractions.